a force of 500N exist between two identical points charge separate by a distance of 40cm.calculate the magnitude of the two point charge with solution.


1
Expert's answer
2022-03-09T12:29:07-0500

Force

"F=\frac{kq_1q_2}{r^2}"

"q_1=q_2=q"

"F=\frac{kq^2}{r^2}"

"q=\sqrt\frac{Fr^2}{k}"

"q=\sqrt\frac{500\times0.16}{9\times10^9}=94.28\times10^{-6}C"

"q_1=q_2=94.28\mu C"